OpenStudy (openstudygirl2):

Evaluate 7 - (-19) - 18 + (-19) +18 will fan and medal for ur help

OpenStudy (openstudygirl2):

helpy

OpenStudy (saksh1998):

7+19-18-19+1 = -10 as -*-=+ +*-=-

OpenStudy (saksh1998):

thank you

OpenStudy (openstudygirl2):

wait

OpenStudy (openstudygirl2):

what do you mean 1? its 18

OpenStudy (openstudygirl2):

@vishal_kothari can u help me @saksh1998 got it wrong

OpenStudy (saksh1998):

sorry misread it 7+19-18-19+18=7

OpenStudy (openstudygirl2):

oh thx

OpenStudy (openstudygirl2):

brb guys i have more evaluating after my 15 minute break

OpenStudy (vishal_kothari):

7 - (-19) - 18 + (-19) +18 = 7+19-18-19+18=7 here 19 got subtracted from 19 becomes 0 and same with 18